Answer items 13 to 15 from the following information:
Professor Dunn is interested in predicting class success on the final examination from scores on the midterm examination. He notes that the correlation between the midterm and final is .80 (a high correlation). Assume that the tests had the following characteristics: mean of midterm = 100; mean of final = 200; s of midterm = 10; s of final = 100; r = .80.

13. Calculate the standard error of estimate.

14. The equation for predicting final examination scores from midterm scores is defined as Ŷ = 3 × X + 10. Terry scored 40 on the midterm. What is Terry’s predicted score on the final?
Answer #1

13)

standard error of estimate =sy*√(1-r2) =100*√(1-.8^2) =60

14)

Terry’s predicted score on the final =3*40+10 =130